OpenStudy (anonymous):

Can someone please check my answer for this? Use synthetic substitution to find g(4) for the function g(x)=x^5-9x^3-10x+8....i got -112...but i don't think its right

OpenStudy (anonymous):

Not correct. Let's go through your work and find the mistake.

OpenStudy (anonymous):

i used that table thing...and substituted the 4 into the g(x) so that the equation was g(4)=x^5-9x^3-10x+8...and then the table looked something like....

OpenStudy (anonymous):

4^5-9*4^3-10*4+8 = 4^5- 9 *12 - 48 256-108 -48 148-48=100

OpenStudy (anonymous):